# The Axiom of Real Determinacy and the Axiom of Real Blackwell Determinacy

### v1

@@ 5. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and $\infty$-Borel codes
The argument is based on the result of Kechris and Woodin [11, Theorem 6.2]. We first introduce the notion of weakly meager sets. A subset $B$ of $\gamma^{\omega}$ is weakly meager if there is an $X\in U_{\pi}$ such that $(\forall b\in X)\,b^{\omega}\cap B$ is meager in the space $b^{\omega}$. Since $b$ is countable, the space $b^{\omega}$ is homeomorphic to the Baire space in most cases. Note that if $B$ is a meager set in the space $\gamma^{\omega}$, then it is weakly meager. A subset $B$ of $\gamma^{\omega}$ is weakly comeager if its complement is weakly meager. Let $I_{\mathrm{wm}}$ be the set of all weakly meager sets.
Sublemma 45.
  1. (1)
    The ideal $I_{\mathrm{wm}}$ is a $\sigma$-ideal on $\gamma^{\omega}$.
  2. (2)
    For any $s\in\gamma^{<\omega}$, the set $s]=\{x\in\gamma^{\omega}\mid s\subseteq x\}$ is not weakly meager.
  3. (3)
    If a subset $B$ of $\gamma^{\omega}$ is not weakly meager, then there is an $s\in\gamma^{<\omega}$ such that $s]\setminus B$ is weakly meager.
  4. (4)
    Let $g$ be a function from $\gamma^{\omega}$ to $\mathrm{Ord}$. Then for any $B$ which is not weakly meager, there is a $B^{\prime}\subseteq B$ which is not weakly meager such that for all $x$ and $y$ in $B^{\prime}$, if $\mathrm{ran}(x)=\mathrm{ran}(y)$, then $g(x)=g(y)$.
Proof.
The first item follows from the $\omega_{1}$-completeness of $U_{\pi}$. The second item follows from the fineness of $U_{\pi}$.
For the third item, suppose $B$ is not weakly meager. Then since $U_{\pi}$ is an ultrafilter, there is an $X\in U_{\pi}$ such that $(\forall b\in X)$ $b^{\omega}\cap B$ is not meager in $b^{\omega}$. Take any $b$ in $X$. Since the space $b^{\omega}$ is homeomorphic to the Baire space, the set $b^{\omega}\cap B$ has the Baire property in $b^{\omega}$. Hence there is an $s_{b}\in b^{<\omega}$ such that $s_{b}]\setminus B$ is meager in $b^{\omega}$. By normality of $U_{\pi}$, there is a $Y\in U_{\pi}$ such that $Y\subseteq X$ and there is an $s\in\gamma^{<\omega}$ such that $s_{b}=s$ for any $b\in Y$. Hence $s]\setminus B$ is weakly meager.
For the last item, let $g$ be such a function and $B$ be not weakly meager. Then there is an $X\in U_{\pi}$ such that $\forall b\in X$, $b^{\omega}\cap B$ is not meager in $b^{\omega}$. Since $b^{\omega}\cap B$ has the Baire property in $b^{\omega}$, there is an $s_{b}\in b^{<\omega}$ such that $s_{b}]\setminus B$ is meager in $b^{\omega}$. By normality of $U_{\pi}$, there are a $Y\subseteq X$ and $s_{0}\in\gamma^{<\omega}$ such that $Y\in U_{\pi}$ and $s_{b}=s_{0}$ for every $b\in Y$.
Take any $b\in Y$. Since $s_{0}]\cap b^{\omega}$ is homeomorphic to the Baire space, by Theorem 3, the meager ideal on the space $s_{0}]\cap b^{\omega}$ is closed under well-ordered unions. Hence there is an $\alpha_{b}$ such that $s_{0}]\cap b^{\omega}\cap g^{-1}(\alpha_{b})$ is not meager in $s_{0}]\cap b^{\omega}$. Since the set $s_{0}]\cap b^{\omega}\cap g^{-1}(\alpha_{b})$ has the Baire property in $s_{0}]\cap b^{\omega}$, there is a $t_{b}\in b^{<\omega}$ such that $t_{b}\supseteq s_{0}$ and $t_{b}]\setminus g^{-1}(\alpha_{b})$ is meager in $b^{\omega}$. By normality of $U_{\pi}$, there are a $Z\in U_{\pi}$ with $Z\subseteq Y$ and a $t_{0}\in b^{<\omega}$ such that $t_{0}\supseteq s_{0}$ and for all $b\in Z$, $t_{b}=t_{0}$. So for all $b\in Z$, $t_{0}]\setminus g^{-1}(\alpha_{b})$ is meager in $b^{\omega}$. Then $B^{\prime}=B\cap[t_{0}]\cap\{x\mid g(x)=\alpha_{\mathrm{ran}(x)}\}$ is the desired one. ∎
Now we prove Lemma 44. Let $G$ be $\mathrm{Coll}(\omega,\gamma)$-generic over $V$. Consider the Boolean algebra $\wp(\gamma^{\omega})/I_{\mathrm{wm}}$. Then it is naturally forcing equivalent to $\mathrm{Coll}(\omega,\gamma)$: In fact, for $s\in\gamma^{<\omega}$, let $i(s)=[s]/I_{\mathrm{wm}}$. Then by the second and third items of Sublemma 45, $i$ is a dense embedding from $\mathrm{Coll}(\omega,\gamma)$ to $\left(\wp(\gamma^{\omega})/I_{\mathrm{wm}}\right)\setminus\{\mathbf{0}\}$. Define $U^{\prime}$ as follows: For a subset $B$ of $\gamma^{\omega}$ in $V$, $B$ is in $U^{\prime}$ if there is a $p\in G$ such that $p]\setminus B$ is weakly meager. By the genericity of $G$ and the third item of Sublemma 45, $U^{\prime}$ is a $V$-ultrafilter on $(\gamma^{\omega})^{V}$ and $U^{\prime}$ contains all the weakly comeager sets. Take an ultrapower $\mathrm{Ult}\bigl(\mathrm{L}(\mathbb{R},S,f_{0},\pi),U^{\prime}\bigr)=\left(\mathrm{L}(\mathbb{R},S,f_{0},\pi)^{(\gamma^{\omega})^{V}}\cap V\right)/U^{\prime}$ and let $j$ be the ultrapower map. (Note that we consider $\mathrm{L}(\mathbb{R},S,f_{0},\pi)$-valued functions in $V$ which are not necessarily in $\mathrm{L}(\mathbb{R},S,f_{0},\pi)$.)
We show that $j$ is the desired map. We first check Loś’s Theorem for this ultrapower. It is enough to show that for any $B\in U^{\prime}$ and any function $F$ in $V$ from $B$ to $\mathrm{L}(\mathbb{R},S,f_{0},\pi)$ such that all the values of $F$ are nonempty, there is a function $f$ on $B$ in $V$ such that $f(x)\in F(x)$ for all $x$ in $B$. Since there is a surjection from $\mathbb{R}\times\mathrm{Ord}$ to $\mathrm{L}(\mathbb{R},S,f_{0},\pi)$, we may assume that the values of $F$ are sets of reals. But then by using uniformization for every relation on the reals by Theorem 19, we obtain the desired $f$.
Next, we check the well-foundedness of $\mathrm{Ult}\bigl(\mathrm{L}(\mathbb{R},S,f_{0},\pi),U^{\prime}\bigr)$. By $\mathsf{DC}$, we know that the ultrapower $\mathrm{Ult}(V,U_{\pi})$ in $V$ is well-founded. Hence it suffices to show the following: For a function $f\colon\wp_{\omega_{1}}(\gamma)\to\mathrm{Ord}$, let $g_{f}\colon\gamma^{\omega}\to\mathrm{Ord}$ be as follows: $g_{f}(x)=f\bigl(\mathrm{ran}(x)\bigr)$.
Sublemma 46.
The function $f]_{U_{\pi}}\mapsto[g_{f}]_{U^{\prime}}$ is an isomorphism from the structure $\bigl(\bigl(\mathrm{Ord}^{(\wp_{\omega_{1}}(\gamma)})\cap V\bigr)/U_{\pi},\in_{U_{\pi}}\bigr)$ to the structure $\bigl(\bigl(\mathrm{Ord}^{(\gamma^{\omega})}\cap V\bigr)/U^{\prime},\in_{U^{\prime}}\bigr)$.
Proof of Sublemma 46.
We first show that if $f_{1}\in_{U_{\pi}}f_{2}$, then $g_{f_{1}}\in_{U^{\prime}}g_{f_{2}}$. Since $f_{1}\in_{U_{\pi}}f_{2}$, there is an $X\in U_{\pi}$ such that for all $b$ in $X$, $f_{1}(b)\in f_{2}(b)$. Fix a $b$ in $X$. Since the set $\{x\in b^{\omega}\mid\mathrm{ran}(x)=b\}\cap b^{\omega}$ is comeager in $b^{\omega}$, the set $\{x\in b^{\omega}\mid f_{1}\bigl(\mathrm{ran}(x)\bigr)\in f_{2}\bigl(\mathrm{ran}(x)\bigr)\}$ is comeager in $b^{\omega}$. Hence for every $b\in X$, the set $\{x\in b^{\omega}\mid g_{f_{1}}(x)\in g_{f_{2}}(x)\bigr)\}$ is comeager in $b^{\omega}$ and the set $\{x\in\gamma^{\omega}\mid g_{f_{1}}(x)\in g_{f_{2}}(x)\}$ is weakly comeager and hence is in $U^{\prime}$. Therefore, $g_{f_{1}}\in_{U^{\prime}}g_{f_{2}}$. In the same way, one can show that if $f_{1}=_{U_{\pi}}f_{2}$, then $g_{f_{1}}=_{U^{\prime}}g_{f_{2}}$.
Next, we show that the map is surjective. Take any function $g\colon\gamma^{\omega}\to\mathrm{Ord}$ in $V$. We show that there is an $f\colon\wp_{\omega_{1}}(\gamma)\to\mathrm{Ord}$ in $V$ such that $g_{f}=_{U^{\prime}}g$. By Sublemma 45 and the genericity of $G$, there is an $Y$ in $U^{\prime}$ such that if $x$ and $y$ are in $Y$ with the same range, then $g(x)=g(y)$. Since $Y$ is in $U^{\prime}$, there is a $p\in G$ such that $p]\setminus Y$ is weakly meager, hence there is an $X$ in $U_{\pi}$ such that for all $b$ in $X$, $([p]\setminus Y)\cap b^{\omega}$ is meager in $b^{\omega}$. Since $\{x\in b^{\omega}\mid\text{ran}(x)=b\}$ is comeager in $b^{\omega}$, $g$ is constant on a comeager set in $p]\cap b^{\omega}$ for each $b\in X$. Let $\alpha_{b}$ be the constant value for each $b\in X$. Now let $f\colon\wp_{\omega_{1}}(\gamma)\to\mathrm{Ord}$ be such that $f(b)=\alpha_{b}$ if $b$ is in $Y$ and $f(b)=0$ otherwise. Then it is easy to check that $g_{f}=_{U^{\prime}}g$, as desired. ∎
We have shown that $j$ is elementary and that the target model of $j$ is well-founded. So we may assume that the target model of $j$ is transitive. Then $j$ is an elementary embedding from $\mathrm{L}(\mathbb{R},S,f_{0},\pi)$ to $\mathrm{L}\bigl(j(\mathbb{R}),j(S),j(f_{0}),j(\pi)\bigr)$. Let $M=\mathrm{L}\bigl(j(\mathbb{R}),j(S),j(f_{0}),j(\pi)\bigr)$. We finally check that all the reals in $V[G]$ are in $M$. Let $x$ be a real in $V[G]$ and $\tau$ be a $\mathbb{P}$-name for a real in $V$ such that $\tau^{G}=x$. We claim that $f_{\tau}]_{U^{\prime}}=x$, where $f_{\tau}$ is the Baire measurable function from $\mathrm{St}\bigl(\mathrm{Coll}(\omega,\gamma)\bigr)$ to the reals induced by $\tau$ from Lemma 23. Notice that the spaces $\mathrm{St}\bigl(\mathrm{Coll}(\omega,\gamma)\bigr)$ and $\gamma^{\omega}$ are isomorphic as Baire spaces, i.e., for some comeager subset $D$ of $\mathrm{St}\bigl(\mathrm{Coll}(\omega,\gamma)\bigr)$, $D$ and $\gamma^{\omega}$ are homeomorhpic. So the notation $f_{\tau}]_{U^{\prime}}$ makes sense, and verifying $f_{\tau}]_{U^{\prime}}=x$ will complete the proof of Lemma 44.
Take any natural number $n$ and set $i=x(n)$ where $i=0$ or $1$. We show that $f_{\tau}]_{U^{\prime}}(n)=i$. We may assume $i=1$. Since $x(n)=1$, there is a $p\in G$ such that $p\Vdash\check{n}\in\tau$. By the definition of $f_{\tau}$, for any $x\in[p]$, $f_{\tau}(x)(n)=1$. Since $p$ is in $G$, by the definition of $U^{\prime}$, the set $\{y\mid f_{\tau}(y)(n)=1\}$ is in $U^{\prime}$. Hence $f_{\tau}]_{U^{\prime}}(n)=1$, as desired. ∎
@@
We go back to the proof of Theorem 43. Let us keep using $M$ to denote the model $\mathrm{L}\bigl(j(\mathbb{R}),j(S),j(f_{0}),j(\pi)\bigr)$.
@@
We now argue that $B_{j[S]}\subseteq B_{j(S)}$ in $M$. Let $b\in\wp_{\omega_{1}}(\mathbb{R}^{V})$ be such that $\pi[b]=S$ in $M$. Let $a\in\wp_{\omega_{1}}(\mathbb{R}^{V})$ be the closure of $b$ under $f_{0}$ in $V[G]$. Then since $a$ is countable in $V[G]$, by Lemma 44, $a$ is also in $M$. Then $j(f_{0})[a]=j(f_{0})\bigl[j[a]\big]=j\bigl[f_{0}[a]\bigr]\subseteq j[a]=a$. So $a$ is closed under $j(f_{0})$ in $M$. Hence $j(S){\upharpoonright}j(\pi)[a]$ is an $\infty$-Borel code and $B_{j(S){\upharpoonright}j(\pi)[a]}\subseteq B_{j(S)}$. Since $j(\pi)[a]=j(\pi)\bigl[j[a]\bigr]=j\bigl[\pi[a]\bigr]\supseteq j\bigl[\pi[b]\bigr]=j[S]$ while $j(\pi)[a]=j\bigl[\pi[a]\bigr]\subseteq j[\pi[\mathbb{R}^{V}]]=j[\gamma]$, $j(S)\upharpoonright j(\pi)[a]=j(S)\upharpoonright j[S]$. Since $j(S)\upharpoonright j[S]$ is the same as $j[S]$ as $\infty$-Borel codes, $B_{j[S]}\subseteq B_{j(S)}$ in $M$ as well, as desired.
@@ 7. Equiconsistency between $\mathsf{AD}_{\mathbb{R}}$ and $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$
We first verify that $\Gamma_{\gamma}=\wp(\mathbb{R})\cap\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$. The incusion $\Gamma_{\gamma}\subseteq\wp(\mathbb{R})\cap\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$ is obvious. We argue for the reverse inclusion. Let $A$ be a set of reals in $\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$. We see that $A$ is in $\Gamma_{\gamma}$. Since $A$ is in $\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$, there is a set of reals $B$ such that $B$ is in $\Gamma_{\gamma}$ and $A$ is ordinal definable from $B$ and a real. Since $B$ is in $\Gamma_{\gamma}$, $\Theta_{B}<\gamma$. But since $A$ is ordinal definable from $B$ and a real, $\Theta_{A}\leq\Theta_{B}$. So $\Theta_{A}<\gamma$ and $A$ is in $\Gamma_{\gamma}$, as desired.

### v2

@@ 5. $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and $\infty$-Borel codes
The argument is the same as that for the result of Kechris and Woodin [11, Theorem 6.2].
In the result of Kechris and Woodin, they assumed $V=\mathrm{L}(\mathbb{R})$ and $\gamma$ is a reliable, and considered an elementary embedding $j\colon\mathrm{L}(\mathbb{R})\to\mathrm{L}\bigl(j(\mathbb{R})\bigr)$ in $V[G]$. In our setting, we assume $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$ and $\mathsf{DC}$ while $\gamma$ is an arbitrary ordinal less than $\Theta$, and consider an elementary embedding $j\colon\mathrm{L}(\mathbb{R},S,f_{0},\pi)\to\mathrm{L}\bigl(j(\mathbb{R}),j(S),j(f_{0}),j(\pi)\bigr)$ in $V[G]$. We do not need to assume that $\gamma$ is reliable because we have uniformization for every relation on the reals by Theorem 19, which ensures that Loś’s Theorem holds for the ultrapower witnessing the desired embedding $j$. ∎
@@
We go back to the proof of Theorem 43. Let $G$ and $j$ be as in Lemma 44. Let $M$ be the model $\mathrm{L}\bigl(j(\mathbb{R}),j(S),j(f_{0}),j(\pi)\bigr)$. So $j\colon\mathrm{L}(\mathbb{R},S,f_{0},\pi)\to M$ is in $V[G]$.
@@
We now argue that $B_{j[S]}\subseteq B_{j(S)}$ in $M$. Let $b\in\wp_{\omega_{1}}(\mathbb{R}^{V})$ be such that $\pi[b]=S$ in $M$. Let $a\in\wp_{\omega_{1}}(\mathbb{R}^{V})$ be the closure of $b$ under $f_{0}$ in $V[G]$. Then since $a$ is countable in $V[G]$, by Lemma 44, $a$ is also in $M$. Then $j(f_{0})[a]=j(f_{0})\bigl[j[a]\big]=j\bigl[f_{0}[a]\bigr]\subseteq j[a]=a$. So $a$ is closed under $j(f_{0})$ in $M$. Hence $j(S){\upharpoonright}j(\pi)[a]$ is an $\infty$-Borel code and $B_{j(S){\upharpoonright}j(\pi)[a]}\subseteq B_{j(S)}$ in $M$. Since $j[S]=j\bigl[\pi[b]\bigr]\subseteq j\bigl[\pi[a]\bigr]=j(\pi)\bigl[j[a]\bigr]=j(\pi)[a]$ while $j(\pi)[a]=j\bigl[\pi[a]\bigr]\subseteq j[\pi[\mathbb{R}^{V}]]=j[\gamma]$, we have $j[S]\subseteq j(\pi)[a]\subseteq j[\gamma]$. Hence $j(S)\upharpoonright j[S]\subseteq j(S)\upharpoonright j(\pi)[a]\subseteq j(S)\upharpoonright j[\gamma]$. However, $j(S)\upharpoonright j[S]=j(S)\upharpoonright j[\gamma]$. Therefore, $j(S)\upharpoonright j[S]=j(S)\upharpoonright j(\pi)[a]$. Since $j(S)\upharpoonright j[S]$ is the same as $j[S]$ as $\infty$-Borel codes, $B_{j[S]}=B_{j(S){\upharpoonright}j(\pi)[a]}\subseteq B_{j(S)}$ in $M$, as desired.
@@ 7. Equiconsistency between $\mathsf{AD}_{\mathbb{R}}$ and $\mathsf{Bl}\text{-}\mathsf{AD}_{\mathbb{R}}$
We first verify that $\Gamma_{\gamma}=\wp(\mathbb{R})\cap\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$. The inclusion $\Gamma_{\gamma}\subseteq\wp(\mathbb{R})\cap\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$ is obvious. We argue for the reverse inclusion. Let $A$ be a set of reals in $\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$. We see that $A$ is in $\Gamma_{\gamma}$. Since $A$ is in $\text{HOD}_{\mathbb{R}\cup\Gamma_{\gamma}}$, there is a set of reals $B$ such that $B$ is in $\Gamma_{\gamma}$ and $A$ is ordinal definable from $B$ and a real. Since $B$ is in $\Gamma_{\gamma}$, $\Theta_{B}<\gamma$. But since $A$ is ordinal definable from $B$ and a real, $\Theta_{A}\leq\Theta_{B}$. So $\Theta_{A}<\gamma$ and $A$ is in $\Gamma_{\gamma}$, as desired.